OpenStudy (anonymous):

Aretha has $0.85 in nickels and dimes. She has 2 more nickels than dimes. How many nickels and dimes does she have? Please help!! and show work

OpenStudy (jamesj):

Let N = # of nickels D = # of dimes We need to write down two equations with those two variables. What is the equation you can write down from this part of the problem: "She has 2 more nickels than dimes"?

OpenStudy (anonymous):

n=d + 2?

OpenStudy (jamesj):

correct. Now what's the other equation: "Aretha has $0.85 in nickels and dimes"

OpenStudy (anonymous):

n + d= 0.85?

OpenStudy (jamesj):

No, n + d is the number of coins.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

.10d + 0.05n = 0.85

OpenStudy (jamesj):

Yes. Now you have two equations with the two variables. Solve them.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (jamesj):

btw, decimals annoy me. Your second equation will be easier if you write it cents: 5N + 10D = 85

OpenStudy (anonymous):

thanks :)